OpenStudy (anonymous):

What is the solution set of x 2 + 5x - 5 = 0?

OpenStudy (anonymous):

use the quadratic formula: - B ± √ B2-4AC x = ———————— 2A In our case, A = 1 B = 5 C = -5 Accordingly, B2 - 4AC = 25 - (-20) = 45 Applying the quadratic formula : -5 ± √ 45 x = ————— 2 Can √ 45 be simplified ? Yes! The prime factorization of 45 is 3•3•5 To be able to remove something from under the radical, there have to be 2 instances of it (because we are taking a square i.e. second root). √ 45 = √ 3•3•5 = ± 3 • √ 5 √ 5 , rounded to 4 decimal digits, is 2.2361 So now we are looking at: x = ( -5 ± 3 • 2.236 ) / 2 Two real solutions: x =(-5+√45)/2=(-5+3√ 5 )/2= 0.854 or: x =(-5-√45)/2=(-5-3√ 5 )/2= -5.854

OpenStudy (anonymous):

dont have any of those as answers

OpenStudy (anonymous):

what answers u have?

OpenStudy (anonymous):